# Dual-Scale Antenna Deployment for Pinching Antenna Systems

### v1

@@ Abstract
A dual-scale deployment (DSD) framework for pinching antenna systems (PASS) is proposed. 1) In the first coarse stage, the pinching antenna (PA) is transferred over a large-scale range at the waveguide level. 2) The refinement stage performs small-scale relocation of the PA with high precision. Four PA deployment protocols are provided in the proposed DSD framework. Then, a practical power consumption model is proposed, based on which the theoretical energy efficiency formulas for PASS are derived. The transmit precoding, PA radiation power, and PA deployment are jointly optimized to maximize the energy efficiency under the provided PA deployment protocols. To solve this non-convex, highly coupled problem, a low-complexity penalty-based alternating optimization algorithm is proposed. Simulation results validate the accuracy of theoretical results and the convergence of the proposed algorithm. It is demonstrated that: 1) PASS delivers about $70\%$ higher energy efficiency than the conventional cell-free architecture and nearly twofold improvement relative to MIMO systems; 2) it is essential to specify the DSD resolution and deployment protocol to achieve the maximum energy efficiency for PASS.
@@ I Introduction
The advent of sixth-generation (6G) wireless systems is anticipated to impose far more stringent requirements on spectral efficiency and full-dimensional coverage[1]. To meet these demands, emerging technologies such as reconfigurable intelligent surfaces [2, 3], fluid antennas [4, 5], and movable antennas [6] have shown promising performance improvements. These flexible antenna technologies advance transceiver architectures and enable environmental reconfigurability, which allows for controllable signal shaping. However, the potential benefits of these technologies are limited as they do not fully mitigate the performance degradation caused by large-scale fading.
Recently, pinching antenna systems (PASS) [7, 8, 9] have gained significant attention for their ability to boost spectral efficiency by enabling near-wire communications. Specifically in PASS, the dielectric waveguide serves both as the transport medium and the antenna rail, while small dielectric particles are pinched onto the waveguide to form pinching antennas (PAs), which act as programmable leakage points for signal radiation. Compared with prior flexible-antenna approaches, PASS directs part of the signal propagation through low-attenuation waveguides and radiates it closer to users, thereby reconfiguring large-scale fading and significantly reducing path loss. This additional spatial degree-of-freedom (DoF) allows for the optimization of PA deployment and power radiation, enabling PA beamforming to shape the emitted field. These capabilities enhance the communication service quality, such as facilitating unobstructed communication links [10] and improving physical-layer security [11, 12]. Therefore, PA technology has drawn tremendous attention from both academia and industry[13], and is expected to play an important role in future wireless communications.
@@ I-A Prior Works
Motivated by the aforementioned favorable characteristics of PASS, substantial research has been conducted to optimize the design of PA power radiation and deployment [18, 19, 17, 14, 15, 16], aiming to fully exploit the additional DoFs introduced by PA and achieve performance gains in PASS. Studies on the PA power radiation typically categorize the methods into general, equal, and proportional power allocation, while PA deployment techniques are generally classified into discrete, continuous, and semi-continuous activation modes.
More specifically, the authors of [14] introduced both equal-power and proportional-power radiation schemes under a physics-based hardware model, formulating a transmit power minimization problem that jointly optimizes transmit and pinching beamforming. Due to its simplicity, the equal-power radiation approach was adopted in [15] and [16], where it was applied to multicast scenarios, with pinching beamforming designed to maximize the multicast rate. In contrast, the general-power scheme offers more flexibility in power allocation across PAs with the potential for higher performance, although its hardware implementation is more complex. For instance, a general-power radiation scheme was proposed in [17], yielding significant transmit power savings compared to the equal-power approach. In terms of PA deployment optimization, a practical strategy involves pre-deploying a large number of PAs and discretely activating a subset to serve communication users. In this regard, the authors of [18] formulated a transmit power minimization problem that optimizes transmit beamforming, pinching beamforming, and the number of active PAs, while ensuring the rate constraint of individual users is met. Additionally, the authors of [19] proposed three practical PA placement schemes, i.e., edge-deployment, center-deployment, and diagonal-deployment. Based on these PA deployment schemes, they investigated the performance of a novel simultaneous wireless information and power transfer system. Furthermore, continuous PA activation methods have been investigated in [17, 14, 15, 16] to maximize system performance by optimizing PA deployment within the waveguide region, subject to the constraint that the distance between any two PAs exceeds a minimum threshold.
@@ I-B Motivations and Contributions
To effectively address these concerns, we propose a dual-scale deployment (DSD) framework that employs a hierarchical design for PA deployment. In the proposed framework, each PA is physically housed in a mobile base that slides on the waveguide, while the PA itself is tunable within the base. This architecture naturally decouples the PA deployment into two-stage operations, ensuring both wide-range repositioning and high-precision refinement. Based on the proposed DSD framework, we design an energy-efficient algorithm for joint optimization of the transmit precoding, PA power radiation, and PA deployment. The main contributions of this paper are summarized as follows:
  • We propose the DSD framework for PA deployment in PASS. In the proposed DSD framework, the mobile base executes large-scale sliding in the coarse stage, while the PA performs small-scale tuning in the refinement stage. We provide four distinct implementation protocols within the DSD framework, namely: “base sliding then PA tuning (STT)”, “base selection then PA activation (STA)”, “base selection and PA tuning (SAT)”, and “base selection and PA activation (SAA)”, along with their respective advantages and limitations.
  • We propose a realistic PASS power consumption model. Based on the proposed model, we derive closed-form theoretical expressions for energy efficiency in the DSD framework. Then, we formulate an energy efficiency maximization problem to jointly optimize the transmit precoding, PA radiation power, and PA deployment under four implementation protocols.
  • We propose a novel low-complexity algorithm to efficiently address the energy efficiency maximization problem. In particular, we solve the PA deployment optimization by decoupling it into two independent subproblems in the DSD framework, i.e., optimizing the coarse-stage base placement and refinement-stage PA deployment. The proposed algorithm also incorporates Lagrangian transformations and block coordinate manifold optimization to tackle the transmit precoding and PA power radiation problems.
  • We provide comprehensive numerical results to validate the accuracy of derived theoretical results and the convergence of the proposed algorithm. Numerical results also demonstrate that: 1) PASS delivers about $70\%$ higher energy efficiency than the conventional cell-free architecture and nearly a twofold improvement relative to MIMO systems; 2) the impact of DSD resolutions on energy efficiency is more pronounced in multi-PA systems; 3) the STT protocol provides over $80\%$ improvement in energy efficiency due to the sliding-tuning gain, while the SAT protocol delivers the highest energy efficiency among the four protocols.
@@ II-B Four Implementation Protocols in the DSD Framework
It can be observed that the challenge of solving problem (11) lies in the coupling between $a_{nk}$ and $\theta_{nk}$, $\forall n$. To solve this, we propose a two-stage optimization method. The insight behind this method is that a small variation of the PA position $\Delta\mathbf{p}_{n}$ induces a negligible change in amplitude term $\alpha_{nk}(\mathbf{p}_{n}+\Delta\mathbf{p}_{n})$, but enables $\theta_{nk}(\mathbf{p}_{n}+\Delta\mathbf{p}_{n})$ to sweep the full range of $\pi,\pi)$. Hence, $\theta_{nk}$ can be tuned by small PA movements with negligible amplitude perturbation, which validates the decoupling approximation and motivates the DSD framework. In this way, we first solve the decoupled problem as
@@
Based on the optimal $a_{nk}^{\mathrm{opt}}$ and $\theta_{nk}^{\mathrm{opt}}$ derived from the problem (12), the proposed DSD strategy carries out coarse- and refinement-stage of the $n$-th PA deployment in a two-stage manner. Particularly, in the coarse stage, the $n$-th PA positioning is adjusted to $\mathbf{p}_{n}^{C}$ to achieve the desired amplitude by solving
@@
Then, in the refinement stage, the $n$-th PA will be further refined with $\Delta\mathbf{p}_{n}^{F}$ to achieve the desired phase by solving
@@
  • STT: Base sliding is carried out by a motorized module over a waveguide-level range for coarse-stage deployment. Fine-tuning PA inside the base is then performed by a piezoelectric module.
  • STA: Base sliding is carried out by a motorized module over a waveguide-level range for coarse-stage deployment. $N_{A}$ PAs are pre-deployed inside the base, where one PA is activated each time slot.
  • SAT: $N_{B}$ bases are pre-deployed along the waveguide, each equipped with a tunable PA. In each time slot, one base is selected, and the PA within that base is finely tuned by a piezoelectric module.
  • SAA: $N_{B}$ bases are pre-deployed along the waveguide, each containing $N_{A}$ PAs. In each time slot, one base is selected, and one PA within that PB is activated.
@@
In Table I, we summarize the response time, pre-deployed PA number, power consumption, and spectral efficiency of these four implementation protocols (a)-(d). Across the four implementation protocols, the PA deployment algorithm designs only differ in the search grid sets in Eqs. (13) and (II-B). We next detail the protocol-specific grid constructions. For simplified mathematical modeling, we assume that the waveguide is deployed along the $x$-axis. This means that the PA deployment designs only require consideration of its $x$-coordinate. Let $x^{\min}$ and $x^{\max}$ denote the minimum and maximum coordinates of the base deployment range along the waveguide, respectively. We use $\delta^{C}$ and $\delta^{F}$ to indicate the position adjustment step sizes, or resolutions, of the sliding base and tuning PA, respectively. Besides, let $L^{F}$ denote the length of PA’s deployment range within each base in the refinement stage. As for the selection or activation representations, it is assumed that $N\cdot N^{F}$ bases and $N\cdot N^{C}$ PAs are pre-deployed, from which $N$ bases or PAs are selected/activated in each time slot. Four practical implementations of the proposed DSD search grid set are provided as follows:
@@ IV-B Penalty-based Alternating Optimization Algorithm
In the proposed DSD framework, the two-stage approach performs two sequential one-dimensional searches with total complexity on the order $\mathcal{O}(G_{C}+G_{F})$, where $G_{C}$ and $G_{F}$ are the search grids number of coarse-stage and refinement-stage, respectively. It substantially reduces the computational complexity compared with prior methods whose joint search [14] has complexity on the order of $\mathcal{O}(G_{C}G_{F})$.
@@
Algorithm 2 Proposed Alternating Optimization Algorithm for Joint BS Transmit Precoding, PA Power Radiation, and PA Positioning.
@@ VI Conclusion
The proposed DSD framework for energy-efficient PA deployment could motivate future research on PASS, such as protocol design and channel estimation. More specifically, designing an appropriate protocol in the DSD framework for PASS is a promising direction for achieving an optimal trade-off among system performance, power consumption, and response time. It could also explicitly account for dual-scale position adjustment speeds together with practical hardware constraints. In addition, dual-scale channel estimation and codebook design call for further investigation to accurately acquire CSI for this hybrid architecture.

### v2

@@ Abstract
A dual-scale deployment (DSD) framework is proposed for pinching antenna systems (PASS), under which four protocols are provided. 1) For the coarse-scale deployment, the pinching antenna (PA) is transferred over a large-scale range at the waveguide level. 2) For the fine-scale deployment, the PA is adjusted with high precision within a small-scale region. By simultaneously optimizing both scales, the proposed DSD framework can unleash the full potential of PA deployment, while maintaining low computational complexity. Based on this framework, we establish a practical power consumption model and derive theoretical energy efficiency expressions for PASS. Then, an energy-efficiency maximization problem is formulated to jointly optimize the transmit precoding, PA radiation power, and dual-scale PA deployment. To solve this non-convex, highly coupled problem, a low-complexity penalty-based alternating optimization algorithm is proposed. Simulation results validate the accuracy of theoretical results and the convergence of the proposed algorithm. It is demonstrated that the proposed DSD framework is highly effective for PASS, delivering about $70\%$ higher energy efficiency than the conventional cell-free architecture and nearly a twofold improvement relative to MIMO systems.
@@ I Introduction
The advent of sixth-generation (6G) wireless systems is anticipated to impose far more stringent requirements on spectral efficiency and full-dimensional coverage[1]. To meet these demands, emerging technologies such as reconfigurable intelligent surfaces [2], fluid antennas [3], and movable antennas [4] have shown promising performance improvements. These flexible antenna technologies advance transceiver architectures and enable environmental reconfigurability, which allows for controllable signal shaping. However, the potential benefits of these technologies are limited as they do not fully mitigate the performance degradation caused by large-scale fading. Recently, pinching antenna system (PASS) [5, 6, 7] has gained significant attention for their ability to boost spectral efficiency by enabling near-wire communications. Specifically in PASS, the dielectric waveguide serves both as the transport medium and the antenna rail, while small dielectric particles are pinched onto the waveguide to form pinching antennas (PAs), which act as programmable leakage points for signal radiation. Compared with prior flexible-antenna approaches, PASS directs part of the signal propagation through low-attenuation waveguides and radiates it closer to users, thereby reconfiguring large-scale fading and significantly reducing path loss. This additional spatial degree-of-freedom (DoF) allows for the optimization of PA deployment and power radiation, enabling PA beamforming to shape the emitted field. These capabilities enhance the communication service quality, such as facilitating unobstructed communication links [8] and improving physical-layer security [9, 10]. Therefore, PASS technology has drawn tremendous attention from both academia and industry[11], and is expected to play an important role in future wireless communications.
@@ I-A Prior Works
Motivated by the aforementioned favorable characteristics of PASS, substantial research has been conducted to optimize the design of PA power radiation and deployment [15, 16, 14, 12, 13], aiming to fully exploit the additional DoFs introduced by PA and achieve performance gains in PASS.
More specifically, the authors of [12] introduced both equal-power and proportional-power radiation schemes under a physics-based hardware model, formulating a transmit power minimization problem that jointly optimizes transmit and pinching beamforming. Due to its simplicity, the equal-power radiation approach was adopted in [13], where it was applied to multicast scenarios, with pinching beamforming designed to maximize the multicast rate. In contrast, the general-power scheme offers more flexibility in power allocation across PAs with the potential for higher performance, although its hardware implementation is more complex. For instance, a general-power radiation scheme was proposed in [14], yielding significant transmit power savings compared to the equal-power approach. In terms of PA deployment optimization, a practical strategy involves pre-deploying a large number of PAs and discretely activating a subset to serve communication users. In this regard, the authors of [15] formulated a transmit power minimization problem that optimizes transmit beamforming, pinching beamforming, and the number of active PAs, while ensuring the rate constraint of individual users is met. Additionally, the authors of [16] proposed three practical PA placement schemes, i.e., edge-deployment, center-deployment, and diagonal-deployment. Furthermore, continuous PA activation methods have been investigated in [14, 12, 13] to maximize system performance by optimizing PA deployment within the waveguide region, subject to the constraint that the distance between any two PAs exceeds a minimum threshold.
@@ I-B Motivations and Contributions
To effectively address these concerns, we propose a dual-scale deployment (DSD) framework that employs a hierarchical design for PA deployment. In the proposed framework, each PA is physically housed in a mobile base that slides on the waveguide, while the PA itself is tunable within the base. This architecture naturally transforms the PA deployment into dual-scale operations, ensuring both wide-range repositioning and high-precision refinement. A similar framework is reflected in the two-stage algorithms adopted in [17, 18, 19], but these works rely on the decoupled optimization of these two scales. For instance, [17] minimized the path loss by sequentially moving the PA to the nearest position to the user before phase tuning. However, such decoupled optimization fails to effectively generalize to multi-user interference scenarios, since it overlooks the spatial coupling between the two scales essential for mitigating inter-user interference. To bridge this gap, we design an energy-efficiency maximization algorithm to jointly optimize the transmit precoding, PA power radiation, and dual-scale PA deployment. The main contributions of this paper are summarized as follows:
  • We propose the DSD framework for PA deployment in PASS. In the proposed DSD framework, the mobile base executes coarse-scale sliding, while the PA performs fine-scale tuning. We provide four implementation protocols within the DSD framework, namely: “base sliding then PA tuning (STT)”, “base selection then PA activation (STA)”, “base selection and PA tuning (SAT)”, and “base selection and PA activation (SAA)”, along with their respective advantages and limitations.
  • We propose a realistic PASS power consumption model. Based on the proposed model, we derive closed-form theoretical expressions for energy efficiency in the DSD framework. Then, we formulate an energy efficiency maximization problem to jointly optimize the transmit precoding, PA radiation power, and dual-scale PA deployment under four implementation protocols.
  • We propose a novel low-complexity algorithm to efficiently address the energy-efficiency maximization problem. In particular, we solve the dual-scale PA deployment optimization into two subproblems in the DSD framework, i.e., optimizing the coarse-scale base placement and fine-scale PA deployment. The proposed algorithm also incorporates Lagrangian transformations and block coordinate manifold optimization to tackle the transmit precoding and PA power radiation problems.
  • We provide numerical results to validate the accuracy of derived theoretical results and the convergence of the proposed algorithm. Numerical results also demonstrate that: 1) the proposed DSD framework is highly effective for PASS, delivering about $70\%$ higher energy efficiency than the conventional cell-free architecture and nearly a twofold improvement relative to MIMO systems; 2) the impact of DSD resolutions on energy efficiency is more pronounced in multi-PA systems; 3) the STT protocol provides over $80\%$ improvement in energy efficiency due to the sliding-tuning gain, while the SAT protocol delivers the highest energy efficiency among the four protocols.
@@ II-B Four Implementation Protocols in the DSD Framework
It can be observed that the challenge of solving problem (11) lies in the coupling between $a_{nk}$ and $\theta_{nk}$, $\forall n$. To solve this, we propose a dual-scale optimization method. The insight behind this method is that a small variation of the PA position $\Delta\mathbf{p}_{n}$ induces a negligible change in amplitude term $\alpha_{nk}(\mathbf{p}_{n}+\Delta\mathbf{p}_{n})$, but enables $\theta_{nk}(\mathbf{p}_{n}+\Delta\mathbf{p}_{n})$ to sweep the full range of $\pi,\pi)$. Hence, $\theta_{nk}$ can be tuned by small PA movements with negligible amplitude perturbation, which validates the decoupling approximation and motivates the DSD framework. In this way, we first solve the decoupled problem as
@@
Based on the optimal $a_{nk}^{\mathrm{opt}}$ and $\theta_{nk}^{\mathrm{opt}}$ derived from the problem (12), the proposed DSD strategy carries out coarse- and refinement-deployment of the $n$-th PA in a dual-scale manner. Particularly, for the coarse-scale deployment, the $n$-th PA positioning is adjusted to $\mathbf{p}_{n}^{C}$ to achieve the desired amplitude by solving
@@
Then, for the fine-scale deployment, the $n$-th PA will be further refined with $\Delta\mathbf{p}_{n}^{F}$ to achieve the desired phase by solving
@@
  • STT: Base sliding is carried out by a motorized module over a waveguide-level range for coarse-scale deployment. Fine-tuning PA inside the base is then performed by a piezoelectric module.
  • STA: Base sliding is carried out by a motorized module over a waveguide-level range for coarse-scale deployment. $N^{F}$ PAs are pre-deployed inside the base, where one PA is activated each time slot.
  • SAT: $N^{C}$ bases are pre-deployed along the waveguide, each equipped with a tunable PA. In each time slot, one base is selected, and the PA within that base is finely tuned by a piezoelectric module.
  • SAA: $N^{C}$ bases are pre-deployed along the waveguide, each containing $N^{F}$ PAs. In each time slot, one base is selected, and one PA within that PB is activated.
@@
In Table I, we summarize the response time, pre-deployed PA number, power consumption, and spectral efficiency of these four implementation protocols (a)-(d). Across the four implementation protocols, the PA deployment algorithm designs only differ in the search grid sets in Eqs. (13) and (II-B). We next detail the protocol-specific grid constructions. For simplified mathematical modeling, we assume that the waveguide is deployed along the $x$-axis. This means that the PA deployment designs only require consideration of its $x$-coordinate. Let $x^{\min}$ and $x^{\max}$ denote the minimum and maximum coordinates of the base deployment range along the waveguide, respectively. We use $\delta^{C}$ and $\delta^{F}$ to indicate the position adjustment step sizes, or resolutions, of the sliding base and tuning PA, respectively. Besides, let $L^{F}$ denote the length of PA’s deployment range within each base for the refinement. As for the selection or activation representations, it is assumed that $N\cdot N^{F}$ bases and $N\cdot N^{C}$ PAs are pre-deployed, from which $N$ bases or PAs are selected/activated in each time slot. Four practical implementations of the proposed DSD search grid set are provided as follows:
@@ IV-B Penalty-based Alternating Optimization Algorithm
In the proposed DSD framework, the dual-scale approach performs two sequential one-dimensional searches with total complexity on the order $\mathcal{O}(G_{C}+G_{F})$, where $G_{C}$ and $G_{F}$ are the search grids number of coarse-scale and fine-scale optimization, respectively. It substantially reduces the computational complexity compared with prior methods whose joint search [12] has complexity on the order of $\mathcal{O}(G_{C}G_{F})$.
@@
Algorithm 2 Proposed Alternating Optimization Algorithm for Joint BS Transmit Precoding, PA Power Radiation, and Dual-Scale PA Positioning.
@@ VI Conclusion
The proposed DSD framework for energy-efficient PA deployment could motivate future research on PASS, such as protocol design and channel estimation. More specifically, designing an appropriate transmission protocol in the DSD framework for PASS is a promising direction for achieving an optimal trade-off among system performance, power consumption, and response time. It could also explicitly account for dual-scale position adjustment speeds together with practical hardware constraints. In addition, dual-scale channel estimation and codebook design call for further investigation to accurately acquire CSI for this hybrid architecture.